OpenStudy (bananas1234):

Solve the equation. If necessary, round your answer to the nearest tenth. 4x^2 = 121 2.75, –2.75 5.5, –5.5 11, –11 30.25, –30.25

OpenStudy (bananas1234):

@Ahsome do you know how to do this one?

OpenStudy (bananas1234):

@rational

OpenStudy (rational):

divide 4 both sides then notice that 4 and 121 are perfect squares

OpenStudy (rational):

4 = 2^2 121 = 11^2

OpenStudy (bananas1234):

I think that makes sense, thanks